OpenStudy (anonymous):

Evaluate the expression |a| - 7 if a = -23. Simplify. |-13| - |5|.

OpenStudy (jhannybean):

\[\sf |-27| - 7 \implies 27 - 7 =~?\]

OpenStudy (anonymous):

*23

OpenStudy (jhannybean):

\[\sf |-13| - |5| = 13 - 5 =~?\]

OpenStudy (jhannybean):

Sorry, 23 - 7 = ? *

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

Lol, no problem.